Calculate the energy required to convert the water into ice at the temperature of - 20 degree Celsius



1
Expert's answer
2021-11-06T01:30:19-0400

Let's assume that we are given 1 kg of water at 20°C.


The specific heat capacity of water is 4186 J/kg.

The Latent Heat of Fusion for ice is 3840 J/kg.

The specific heat of ice is 2100 J/kg.


Q = m x C x [20°C - 0°C] + m x LF + m x Cice[0° - (-20°C)]


Q = (1.0 kg x 4186 J/kg. C° x 20°C) + (1.0 kg x 3840 J/kg) + (1.0 kg x 2100 J/kg.C° x 20°C) = 83720 + 3840 + 42 000 = 129.6 kJ.